# Multi-fidelity Bayesian Optimization: A Review

### v2

@@ 4.4 Bayesian hierarchical model
Given two computational models with different fidelity for predicting an objective function, and the corresponding predictions $f_{1}({\bf x})=f_{\text{L}}({\bf x})$ and $f_{2}({\bf x})=f_{\text{H}}({\bf x})$, Qian and Wu [97] proposed a Bayesian hierarchical model for approximating the prediction of $f_{\text{2}}({\bf x})$. Using the KOH model, the Bayesian hierarchical model adds a measurement error to the hybrid additive/multiplicative form in Eq. 2 for $w=1$ and $c\neq 0$, such that
@@ 4.5.1 Deep Gaussian processes
A deep Gaussian process (DGP) describes the mapping between the input variables and the output using a composition of GPs [194]. Figure 7 shows a general architecture of DGP with $L$ hidden layers in which a GP relates two consecutive layers. The architecture in Fig. 7 is mathematically described as follows:
@@ 4.5.2 Nonlinear auto-regressive model
Based on the KOH model and an output-output mapping technique, Perdikaris et al. [23] proposed the nonlinear auto-regressive model for learning complex nonlinear correlations between models of variable fidelity. Let $h_{t}(\cdot):\mathbb{R}^{d+1}\mapsto\mathbb{R}$ be a mapping that relates the vector of input and output of fidelity $(t-1)$ to the output of fidelity $t$. Let ${\bf z}_{t-1}({\bf x})=\left[{\bf x}^{\intercal},\hat{f}_{t-1}({\bf x})\right]^{\intercal}$ be the input vector of the surrogate predicting the output of fidelity $t$, where $\hat{f}_{t-1}({\bf x})$ is the GP posterior corresponding to fidelity $(t-1)$. The nonlinear auto-regressive model reads [23]
@@
where $\delta_{1}(\cdot)\sim\mathcal{GP}\left(0,\kappa_{\delta,1}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,1})\right)$ and the condition in Eq. 22c is to describe the Markov property in Eq. 6c.
In a probabilistic framework, the mapping $h_{t}(\cdot)$ in Eq. 22b is modeled as a GP with zero mean and covariance function $\kappa_{\text{h},t}(\cdot,\cdot|{\boldsymbol{\phi}}_{\text{h},t})$, such that
@@
The GP surrogate for the model in Eq. 22b is a DGP when $\delta_{1}(\bf x)$ and $h_{t}(\bf x)$ are GPs. However, the number of hyperparameters for this DGP is much smaller than that for the full DPG described in Section 4.5.1 [23].
For $t>2$, we cannot analytically marginalize the likelihood function described in Eq. 25. As a result, the statistical estimates of $f_{t}$ at an unseen input vector can be obtained via the Monte-Carlo integration.
@@ 4.5.3 Deep multi-fidelity Gaussian process
Using the DGP modeling framework, Cutajar et al. [114] proposed a multi-fidelity deep Gaussian process (MF DGP), with each layer of the DGP corresponding to a fidelity level. Specifically, layers $2,\dots,L$ of the DGP in Fig. 7 correspond to the surrogate models for fidelities $11,\dots,T-$ 1, respectively. This yields the following MF DGP marginal likelihood for the surrogate that approximates the highest-fidelity output:
@@
Since the joint GP prior $p({\bf F}_{t},{\bf U}_{t})$ and the prior $p({\bf U}_{t})$ are Gaussians, the conditional $p({\bf F}_{t}|{\bf U}_{t})$ is also a Gaussian. Thus, both terms of the variational joint posterior in Eq. 26 are Gaussians. This enables marginalizing ${\bf U}_{t}$ from each fidelity level to obtain an analytical form of the variational marginal $q({\bf F}_{t})$, as
@@ 4.6 Input-augmentation multi-fidelity Gaussian processes
The GP-based MF surrogates we have seen so far are built in the space of input variables, or an augmented space of input and output variables. An alternative approach, as briefly described in Section 2.2, constructs GP-based MF surrogates as functions of both input and fidelity variables. We call this approach input-augmentation GP-based MF modeling, and which has been used in BO predominantly. Once an input-augmentation MF surrogate has been trained, the highest-fidelity output at any set of input variables can be predicted by setting the fidelity-level variables at their highest values.
Depending on how the fidelity variables are described, we classify the input-augmentation methods into two groups. The first group relies on continuous approximations of the fidelity variables (Section 4.6.1), while the second group treats them as categorical variables (Section 4.6.2).
@@ 4.6.2 Use of non-continuous covariance functions
where $\varphi(t_{i})=[z_{i,0},\dots,z_{i,l_{i}}]^{\intercal}$ defines the hypersphere mapping, and $\sigma_{\text{t},i}$ represents the cross-correlation between categories of $t_{i}$ and $t_{i}^{\prime}$. The determination of $[z_{i,0},\dots,z_{i,l_{i}}]^{\intercal}$ can be found in Zhou et al. [197] and Pelamatti et al. [198].
Roustant et al. [199] used a so-called compound symmetry covariance function for $\kappa(t_{i},t_{i}^{\prime}|{\boldsymbol{\phi}}_{\text{t},i})$ in Eq. 32b by assuming a common correlation for all categories. Accordingly, $\kappa(t_{i},t_{i}^{\prime}|{\boldsymbol{\phi}}_{t,i})$ reads the following parsimonious form:
@@
Instead of relying on a point-wise product to define $\kappa_{\text{g}}$ as in Eq. 31, Poloczek et al. [201] used a sum of covariance functions to define $\kappa_{\text{g}}$ when maximizing an objective function $f({\bf x})=g(t,{\bf x})-\delta_{t}({\bf x})$. Here $t\in\mathcal{T}=\{t_{1},\dots,t_{T}\}$ is a categorical variable indicating the noisy information source, and therefore can be seen as a fidelity variable. $g(t,{\bf x})$ is the noisy value of $f({\bf x})$ when observing $t$ at ${\bf x}$, thus $g(t_{T},{\bf x})=f({\bf x})$ is the noise-free observation. $\delta_{t}({\bf x})$ is the discrepancy function defined as a GP independent of $f$ such that $\delta_{t}(\cdot)\sim\mathcal{GP}(m_{\delta,t}(\cdot),\kappa_{\delta,t}\left(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,t}\right))$. The objective function is $f(\cdot)\sim\mathcal{GP}(m_{\text{f}}(\cdot),\kappa_{\text{x}}\left(\cdot,\cdot|{\boldsymbol{\phi}}_{\text{x}}\right))$. Thus, $\kappa_{\text{g}}\left((t,{\bf x}),(t^{\prime},{\bf x}^{\prime})|{\boldsymbol{\phi}}_{\text{g}}\right)$ can be derived from $g(t,{\bf x})=f({\bf x})+\delta_{t}({\bf x})$, as
@@ 5 Acquisition functions
An acquisition function formulated in each iteration of BO maps one of our preferences for the next design point to each point in the design variable space. Our preferences, under incomplete knowledge about the objective function, often include an improvement in the objective function value, a gain in information on the true minimizer, and a gain in information on the true minimum. Maximizing the acquisition function, therefore, guides BO toward a new design point at which our preference achieves the highest score. In this section, we first review notable acquisition functions of the generic BO (Section 5.1). We then describe several ways of modifying these acquisition functions for use of MF BO (Section 5.2), followed by a brief discussion on the portfolio of different design points when maximizing a set of multiple acquisition functions (Section 5.3). We finally discuss optimization algorithms used to maximize the acquisition functions (Section 5.4) and recommend open-source software for BO implementation (Section 5.5).
@@ 5.1 Acquisition functions for Bayesian optimization
Two classes of acquisition functions of the generic BO include one-step look-ahead (myopic) and multi-step look-ahead (or non-myopic). One-step look-ahead acquisition functions, which are predominantly used in BO literature, select a new design point using a utility measure without forecasting the potential impact of all future selections beyond the immediate next selection. They can be further classified into improvement-based (Section 5.1.1), optimistic (Section 5.1.2), and information-based (Section 5.1.3) acquisition functions. In contrast, multi-step look-ahead acquisition functions (Section 5.1.4) consider the impact of future selections on the decision of next design points by adding a new term to one-step look-ahead acquisition functions. This aims to mitigate a limitation of one-step look-ahead acquisition functions that they are deemed to prefer exploitation over exploration [202].
Assume that we are at iteration $k$ of BO and wish to select a new design point ${\bf x}^{k}$ by maximizing an acquisition function $\alpha({\bf x})$. What we currently know to formulate $\alpha({\bf x})$, for example, in 9 of Algorithm 1 are the GP posterior $\hat{f}_{\text{H}}^{k}({\bf x})$ constructed from the available (HF) data $\mathcal{D}^{k-1}$ and, under noise-free observations, the best solution $\{{\bf x}_{\min},f_{\min}\}$ we found in $\mathcal{D}^{k-1}$.
@@ 5.1.3 Information-based acquisition functions
1:Input: $f_{\text{H}}(\cdot)$, $\mathcal{X}$, $K$, $N$;
2:Generate $N$ samples of ${\bf x}$;
@@
A pseudo-code for TS is given in Algorithm 2. We see that if the posterior $\hat{f}_{\text{H}}^{k}({\bf x})$ is highly uncertain, TS tends to sample many different posterior functions $q_{\text{H}}^{k}({\bf x})$ in 10, which turns out that TS favors exploration. As the uncertainty reduces, the algorithm starts exploiting the knowledge about the true objective function. Thus, TS handles the exploitation-exploration trade-off in a natural way. In fact, the exploitation ability of TS is inferior because of its inherent randomness [212].
Theoretical results allow for establishing the connection between TS and GP-UCB via the translation from regret bounds developed for GP-UCB into Bayesian regret bounds developed for TS [216]. It is also possible to sample multiple functions $q_{\text{H}}^{k}({\bf x})$ (10 of Algorithm 2) in parallel to select a batch of new design points [217]. Nevertheless, TS may address the following four problem features inadequately: problems that do not require exploration, problems that do not require exploitation, problems that are time-sensitive, and problems that require careful assessment of information gain [218].
@@
where ``"``"``-"` ` - " reformulates the minimization of entropy to the maximization of the corresponding acquisition function, $p(f_{\text{H}}|{\bf x},\mathcal{D}^{k-1})$ denotes the PDF of $f_{\text{H}}$ conditioned on $\bf x$ and $\mathcal{D}^{k-1}$, and $\mathbb{E}_{f_{\text{H}}}\left[\mathbb{H}({\bf x}^{\star}|{\bf x},f_{\text{H}},\mathcal{D}^{k-1})\right]$ is the expected entropy of ${\bf x}^{\star}$ given that the new design point and the corresponding value of the objective function has been added to $\mathcal{D}^{k-1}$.
Since computing the entropy for a continuous ${\bf x}^{\star}$ is generally analytically intractable, discretizing Eq. 45 makes the computation tractable. While IAGO and ES share the same acquisition function form, they differ in their ways of discretizing Eq. 45.
Predictive entropy search (PES) [220] is another information-based acquisition function derived from $\lambda(\mathcal{D}^{k})$. PES arises from an important observation that the maximizer of $-\mathbb{E}_{f_{\text{H}}}\left[\mathbb{H}({\bf x}^{\star}|{\bf x},f_{\text{H}},\mathcal{D}^{k-1})\right]$ is identical to that of $\mathbb{H}({\bf x}^{\star}|\mathcal{D}^{k-1})-\mathbb{E}_{f_{\text{H}}}\left[\mathbb{H}({\bf x}^{\star}|{\bf x},f_{\text{H}},\mathcal{D}^{k-1})\right]$. This is due to the fact that $\mathbb{H}({\bf x}^{\star}|\mathcal{D}^{k-1})$ does not depend on ${\bf x}^{k}$. By further leveraging the mutual information $\mathcal{I}({\bf x}^{\star};f_{\text{H}})$ between ${\bf x}^{\star}$ and $f_{\text{H}}$, we obtain the following relation [221]:
@@
where $\mathbb{E}_{{\bf x}^{\star}}\left[\mathbb{H}(f_{\text{H}}|{\bf x}^{\star},{\bf x},\mathcal{D}^{k-1})\right]$ is the expected entropy of $f_{\text{H}}$ given ${\bf x}$, ${\bf x}^{\star}$, and $\mathcal{D}^{k-1}$. Based on Eq. 46, PES makes use of the following acquisition function:
@@
The calculation of PES is much easier than that of ES because the first and second terms of $\alpha({\bf x})$ in Eq. 47 only require evaluating the entropy of univariate Gaussians. Nevertheless, the calculation of the second term is still intricate as it necessitates marginalization over the multivariate posterior $p({\bf x}^{\star}|\mathcal{D}^{k-1})$ which can be approximated by a finite set of its samples generated by TS.
Rather than using the entropy of unknown minimizer ${\bf x}^{\star}$, maximum-value entropy search (MES) [222, 223] formulates its acquisition function using the entropy of unknown minimum $f_{\text{H}}({\bf x}^{\star})$. In fact, MES modifies PES in Eq. 47 by replacing the minimizer ${\bf x}^{\star}$ of the second term with the minimum $f^{\star}=f_{\text{H}}({\bf x}^{\star})$. Accordingly, MES reads

### v3

@@ 4.4 Bayesian hierarchical model
Given two computational models with different fidelity for predicting an objective function, and the corresponding predictions $f_{1}({\bf x})=f_{\text{L}}({\bf x})$ and $f_{2}({\bf x})=f_{\text{H}}({\bf x})$, Qian and Wu [97] proposed a Bayesian hierarchical model for approximating the prediction of $f_{\text{2}}({\bf x})$. Using the KOH model, the Bayesian hierarchical model adds a measurement error to the hybrid additive/multiplicative form in Eq.˜2 for $w=1$ and $c\neq 0$, such that
@@ 4.5.1 Deep Gaussian processes
A deep Gaussian process (DGP) describes the mapping between the input variables and the output using a composition of GPs [194]. Figure˜7 shows a general architecture of DGP with $L$ hidden layers in which a GP relates two consecutive layers. The architecture in Fig.˜7 is mathematically described as follows:
@@ 4.5.2 Nonlinear auto-regressive model
Based on the KOH model and an output-output mapping technique, Perdikaris et al. [23] proposed the nonlinear auto-regressive model for learning complex nonlinear correlations between models of variable fidelity. Let $h_{t}(\cdot):\mathbb{R}^{d+1}\mapsto\mathbb{R}$ be a mapping that relates the vector of input and output of fidelity (t1)(t-1) to the output of fidelity $t$. Let ${\bf z}_{t-1}({\bf x})=\left[{\bf x}^{\intercal},\hat{f}_{t-1}({\bf x})\right]^{\intercal}$ be the input vector of the surrogate predicting the output of fidelity $t$, where $\hat{f}_{t-1}({\bf x})$ is the GP posterior corresponding to fidelity (t1)(t-1). The nonlinear auto-regressive model reads [23]
@@
where $\delta_{1}(\cdot)\sim\mathcal{GP}\left(0,\kappa_{\delta,1}(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,1})\right)$ and the condition in Eq.˜22c is to describe the Markov property in Eq.˜6c.
In a probabilistic framework, the mapping $h_{t}(\cdot)$ in Eq.˜22b is modeled as a GP with zero mean and covariance function $\kappa_{\text{h},t}(\cdot,\cdot|{\boldsymbol{\phi}}_{\text{h},t})$, such that
@@
The GP surrogate for the model in Eq.˜22b is a DGP when $\delta_{1}(\bf x)$ and $h_{t}(\bf x)$ are GPs. However, the number of hyperparameters for this DGP is much smaller than that for the full DPG described in Section˜4.5.1 [23].
For $t>2$, we cannot analytically marginalize the likelihood function described in Eq.˜25. As a result, the statistical estimates of $f_{t}$ at an unseen input vector can be obtained via the Monte-Carlo integration.
@@ 4.5.3 Deep multi-fidelity Gaussian process
Using the DGP modeling framework, Cutajar et al. [114] proposed a multi-fidelity deep Gaussian process (MF DGP), with each layer of the DGP corresponding to a fidelity level. Specifically, layers $2,\dots,L$ of the DGP in Fig.˜7 correspond to the surrogate models for fidelities $1,\dots,T-1$, respectively. This yields the following MF DGP marginal likelihood for the surrogate that approximates the highest-fidelity output:
@@
Since the joint GP prior $p({\bf F}_{t},{\bf U}_{t})$ and the prior $p({\bf U}_{t})$ are Gaussians, the conditional $p({\bf F}_{t}|{\bf U}_{t})$ is also a Gaussian. Thus, both terms of the variational joint posterior in Eq.˜26 are Gaussians. This enables marginalizing ${\bf U}_{t}$ from each fidelity level to obtain an analytical form of the variational marginal $q({\bf F}_{t})$, as
@@ 4.6 Input-augmentation multi-fidelity Gaussian processes
The GP-based MF surrogates we have seen so far are built in the space of input variables, or an augmented space of input and output variables. An alternative approach, as briefly described in Section˜2.2, constructs GP-based MF surrogates as functions of both input and fidelity variables. We call this approach input-augmentation GP-based MF modeling, and which has been used in BO predominantly. Once an input-augmentation MF surrogate has been trained, the highest-fidelity output at any set of input variables can be predicted by setting the fidelity-level variables at their highest values.
Depending on how the fidelity variables are described, we classify the input-augmentation methods into two groups. The first group relies on continuous approximations of the fidelity variables (Section˜4.6.1), while the second group treats them as categorical variables (Section˜4.6.2).
@@ 4.6.2 Use of non-continuous covariance functions
where $\varphi(t_{i})=[z_{i,0},\dots,z_{i,l_{i}}]^{\intercal}$ defines the hypersphere mapping, and $\sigma_{\text{t},i}$ represents the cross-correlation between categories of $t_{i}$ and $t_{i}^{\prime}$. The determination of $z_{i,0},\dots,z_{i,l_{i}}]^{\intercal}$ can be found in Zhou et al. [197] and Pelamatti et al. [198].
Roustant et al. [199] used a so-called compound symmetry covariance function for $\kappa(t_{i},t_{i}^{\prime}|{\boldsymbol{\phi}}_{\text{t},i})$ in Eq.˜32b by assuming a common correlation for all categories. Accordingly, $\kappa(t_{i},t_{i}^{\prime}|{\boldsymbol{\phi}}_{t,i})$ reads the following parsimonious form:
@@
Instead of relying on a point-wise product to define $\kappa_{\text{g}}$ as in Eq.˜31, Poloczek et al. [201] used a sum of covariance functions to define $\kappa_{\text{g}}$ when maximizing an objective function $f({\bf x})=g(t,{\bf x})-\delta_{t}({\bf x})$. Here $t\in\mathcal{T}=\{t_{1},\dots,t_{T}\}$ is a categorical variable indicating the noisy information source, and therefore can be seen as a fidelity variable. $g(t,{\bf x})$ is the noisy value of $f({\bf x})$ when observing $t$ at $\bf x$}, thus $g(t_{T},{\bf x})=f({\bf x})$ is the noise-free observation. $\delta_{t}({\bf x})$ is the discrepancy function defined as a GP independent of $f$ such that $\delta_{t}(\cdot)\sim\mathcal{GP}(m_{\delta,t}(\cdot),\kappa_{\delta,t}\left(\cdot,\cdot|{\boldsymbol{\phi}}_{\delta,t}\right))$. The objective function is $f(\cdot)\sim\mathcal{GP}(m_{\text{f}}(\cdot),\kappa_{\text{x}}\left(\cdot,\cdot|{\boldsymbol{\phi}}_{\text{x}}\right))$. Thus, $\kappa_{\text{g}}\left((t,{\bf x}),(t^{\prime},{\bf x}^{\prime})|{\boldsymbol{\phi}}_{\text{g}}\right)$ can be derived from $g(t,{\bf x})=f({\bf x})+\delta_{t}({\bf x})$, as
@@ 5 Acquisition functions
An acquisition function formulated in each iteration of BO maps one of our preferences for the next design point to each point in the design variable space. Our preferences, under incomplete knowledge about the objective function, often include an improvement in the objective function value, a gain in information on the true minimizer, and a gain in information on the true minimum. Maximizing the acquisition function, therefore, guides BO toward a new design point at which our preference achieves the highest score. In this section, we first review notable acquisition functions of the generic BO (Section˜5.1). We then describe several ways of modifying these acquisition functions for use of MF BO (Section˜5.2), followed by a brief discussion on the portfolio of different design points when maximizing a set of multiple acquisition functions (Section˜5.3). We finally discuss optimization algorithms used to maximize the acquisition functions (Section˜5.4) and recommend open-source software for BO implementation (Section˜5.5).
@@ 5.1 Acquisition functions for Bayesian optimization
Two classes of acquisition functions of the generic BO include one-step look-ahead (myopic) and multi-step look-ahead (or non-myopic). One-step look-ahead acquisition functions, which are predominantly used in BO literature, select a new design point using a utility measure without forecasting the potential impact of all future selections beyond the immediate next selection. They can be further classified into improvement-based (Section˜5.1.1), optimistic (Section˜5.1.2), and information-based (Section˜5.1.3) acquisition functions. In contrast, multi-step look-ahead acquisition functions (Section˜5.1.4) consider the impact of future selections on the decision of next design points by adding a new term to one-step look-ahead acquisition functions. This aims to mitigate a limitation of one-step look-ahead acquisition functions that they are deemed to prefer exploitation over exploration [202].
Assume that we are at iteration $k$ of BO and wish to select a new design point ${\bf x}^{k}$ by maximizing an acquisition function $\alpha({\bf x})$. What we currently know to formulate $\alpha({\bf x})$, for example, in Line 9 of Algorithm˜1 are the GP posterior $\hat{f}_{\text{H}}^{k}({\bf x})$ constructed from the available (HF) data $\mathcal{D}^{k-1}$ and, under noise-free observations, the best solution $\{{\bf x}_{\min},f_{\min}\}$ we found in $\mathcal{D}^{k-1}$.
@@ 5.1.3 Information-based acquisition functions
1:Input: $f_{\text{H}}(\cdot)$, $\mathcal{X}$, $K$, NN;
2:Generate NN samples of $\bf x$};
@@
A pseudo-code for TS is given in Algorithm˜2. We see that if the posterior $\hat{f}_{\text{H}}^{k}({\bf x})$ is highly uncertain, TS tends to sample many different posterior functions $q_{\text{H}}^{k}({\bf x})$ in Line 10, which turns out that TS favors exploration. As the uncertainty reduces, the algorithm starts exploiting the knowledge about the true objective function. Thus, TS handles the exploitation-exploration trade-off in a natural way. In fact, the exploitation ability of TS is inferior because of its inherent randomness [212].
Theoretical results allow for establishing the connection between TS and GP-UCB via the translation from regret bounds developed for GP-UCB into Bayesian regret bounds developed for TS [216]. It is also possible to sample multiple functions $q_{\text{H}}^{k}({\bf x})$ (Line 10 of Algorithm˜2) in parallel to select a batch of new design points [217]. Nevertheless, TS may address the following four problem features inadequately: problems that do not require exploration, problems that do not require exploitation, problems that are time-sensitive, and problems that require careful assessment of information gain [218].
@@
where ``-" reformulates the minimization of entropy to the maximization of the corresponding acquisition function, $p(f_{\text{H}}|{\bf x},\mathcal{D}^{k-1})$ denotes the PDF of $f_{\text{H}}$ conditioned on $\bf x$ and $\mathcal{D}^{k-1}$, and $\mathbb{E}_{f_{\text{H}}}\left[\mathbb{H}({\bf x}^{\star}|{\bf x},f_{\text{H}},\mathcal{D}^{k-1})\right]$ is the expected entropy of $\bf x$$^{\star}$ given that the new design point and the corresponding value of the objective function has been added to $\mathcal{D}^{k-1}$.
Since computing the entropy for a continuous $\bf x$$^{\star}$ is generally analytically intractable, discretizing Eq.˜45 makes the computation tractable. While IAGO and ES share the same acquisition function form, they differ in their ways of discretizing Eq.˜45.
Predictive entropy search (PES) [220] is another information-based acquisition function derived from $\lambda(\mathcal{D}^{k})$. PES arises from an important observation that the maximizer of $\mathbb{E}_{f_{\text{H}}}\left[\mathbb{H}({\bf x}^{\star}|{\bf x},f_{\text{H}},\mathcal{D}^{k-1})\right]$ is identical to that of $\mathbb{H}({\bf x}^{\star}|\mathcal{D}^{k-1})-\mathbb{E}_{f_{\text{H}}}\left[\mathbb{H}({\bf x}^{\star}|{\bf x},f_{\text{H}},\mathcal{D}^{k-1})\right]$. This is due to the fact that $\mathbb{H}({\bf x}^{\star}|\mathcal{D}^{k-1})$ does not depend on ${\bf x}^{k}$. By further leveraging the mutual information $\mathcal{I}({\bf x}^{\star};f_{\text{H}})$ between $\bf x$$^{\star}$ and $f_{\text{H}}$, we obtain the following relation [221]:
@@
where $\mathbb{E}_{{\bf x}^{\star}}\left[\mathbb{H}(f_{\text{H}}|{\bf x}^{\star},{\bf x},\mathcal{D}^{k-1})\right]$ is the expected entropy of $f_{\text{H}}$ given $\bf x$}, $\bf x$$^{\star}$, and $\mathcal{D}^{k-1}$. Based on Eq.˜46, PES makes use of the following acquisition function:
@@
The calculation of PES is much easier than that of ES because the first and second terms of $\alpha({\bf x})$ in Eq.˜47 only require evaluating the entropy of univariate Gaussians. Nevertheless, the calculation of the second term is still intricate as it necessitates marginalization over the multivariate posterior $p({\bf x}^{\star}|\mathcal{D}^{k-1})$ which can be approximated by a finite set of its samples generated by TS.
Rather than using the entropy of unknown minimizer $\bf x$$^{\star}$, maximum-value entropy search (MES) [222, 223] formulates its acquisition function using the entropy of unknown minimum $f_{\text{H}}({\bf x}^{\star})$. In fact, MES modifies PES in Eq.˜47 by replacing the minimizer $\bf x$$^{\star}$ of the second term with the minimum $f^{\star}=f_{\text{H}}({\bf x}^{\star})$. Accordingly, MES reads